OpenStudy (anonymous):

Mrs. Jones Algebra 2 class scored very well on yesterday’s quiz. With one exception, everyone received an A. Within how many standard deviations from the mean do all the quiz grades fall? 91, 92, 94, 88, 96, 99, 91, 93, 94, 97, 95, 97 A. 2 <--------Double check please B. 1 C. 3 D. 4

OpenStudy (anonymous):

Highest Value = 99 Lowest Value = 88 Number of observations = 12 First calculate the mean and standard deviation, Mean = 93.9 , SD = 3.12 (From analysis) (99-93.9)/3.12 = 1.6 (88 - 93.9)/3.12 = -1.89 all the quiz grades fall within 2 SD from the mean.

OpenStudy (anonymous):

So it A going to be the right answer? I'm a-little confused...